OpenStudy (anonymous):

what are the steps to solve this problem? 0.000027 / 0.000009 explain how to use scientific notation to solve the problem. describe how to divide numbers written in scientific notation. give final answer in simplest form.

OpenStudy (anonymous):

do you know how to write \(0.000027\) in scientific notation? it will make it easy

OpenStudy (anonymous):

yea its 2.7 * 10^-5

OpenStudy (anonymous):

ok do the same with the bottom

OpenStudy (anonymous):

then you get \[\frac{2.7\times 10^{-5}}{9\times 10^{-6}}\] i think

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then \(2.7\div 9=.3\) and \(-5-(-6)=-5+6=1\) giving a final answer of \[.3\times 10^1=3\]

OpenStudy (anonymous):

now that is actually not how i would solve it

OpenStudy (anonymous):

how would you solve it then?

OpenStudy (anonymous):

\[\frac{0.000027 }{ 0.000009}\] i would move the decimal over six places top and bottom, get \[\frac{27}{9}=3\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you for helping me

OpenStudy (anonymous):

yw